In simplifying expressions what is the rule for Adding together a large negative
number an a small positive number?
for example: -15+4

Answer by terbear1963(43)

THE RULES FOR SIGNED NUMBERS ARE EASY!!! JUST REMEMBER THESE RULES : (1) IF THE SIGNS ARE ALIKE, ADD AND KEEP THE SIGN. EX : -4 + -3 = -7 AND (+3) + (+4)= (+7) //// (2) IF THE SIGNS ARE UNLIKE, SUBTRACT AND TAKE THE SIGN OF THE LARGER NUMBER. EX : (-15)+(+4)= -11 AND (-20)+(+35)=(+15) YOU DON'T ALWAYS HAVE TO USE THE ( ) AND THE + SIGN, BUT, IT MAKES IT CLEARER WHEN USING POSITIVE AND NEGATIVE NUMBERS!!! I HOPE THIS HELPS?! E MAIL : cats7051@yahoo.com YOU CAN E MAIL ME IF YOU HAVE ANY MORE ?'S YOUR MATH FRIEND, TERRY :)